The relative permittivity of a die electric is 2.25.(a) What is its permittivity (b) What will be the capacitance of this electric is inserted into a 7.8 air capacitor?



1
Expert's answer
2022-04-21T13:01:26-0400

a)

"\varepsilon _r=\frac{\varepsilon_0}{\varepsilon_m}=3.93~\frac{pF}m,"

b)

"C'=\frac C{7.8}=0.13C."